OpenStudy (anonymous):

log5 sqrt 625 -- evaluate to simplest form

OpenStudy (anonymous):

\[\log_5(625)=\log_5(5^4)=4\]

OpenStudy (anonymous):

you are trying to solve \[5^x=625\] and if you recognize that \[5^4=625\] then the answer is 4

OpenStudy (anonymous):

I think the answer is 2 because it says sqrt 625 which is 25 = 5^2

jhonyy9 (jhonyy9):

because this will be 4log_5(5) this is equal 4

jhonyy9 (jhonyy9):

log_5(5) =1